Each of the four engines can produce a power of 56 MW (megaWatts). Again, assuming no air resistance, how long would it take the plane to reach its cruising speed and altitude? Assume the mass of the plane is constant. Give an answer in minutes and to one decimal place of accuracy.


1
Expert's answer
2021-06-10T16:42:05-0400
"Pt=mgh+0.5mv^2\\4\cdot56\cdot10^6(60\cdot t)=550\cdot10^6(9.8)(13100)\\+0.5(550\cdot10^6)(305)^2\\t=7000\ min"